# Interpreting DESI 2024 BAO: late-time dynamical dark energy or a local effect?

### v1

@@ Abstract
We perform fits to DESI, CMB and supernova data to understand the physical origin of the DESI hint for dynamical dark energy. We find that the linear parametrization of the equation of state $w$ may guide to misleading interpretations, such as the hint for a phantom Universe, which are not preferred by the data. Instead, physical quintessence models fit the data well. Model-independently, present observations prefer deviations from the constant dark energy, $w=-1$, only at very low redshifts, $z<\mathcal{O}(0.1)$. We find that this result is driven by low-$z$ supernova data. Therefore, either the fundamental properties of our Universe, characterised by the equation of state $w$ and the Hubble parameter $H$, underwent dramatic changes very recently or, alternatively, we do not fully understand the systematics of our local Universe in a radius of about $300\,h^{-1}\rm Mpc$.
@@
Introduction – We have come to a turning point where the standard $\Lambda$CDM model, previously regarded as robust, may have to be reconsidered. Taking into account recent measurements of baryon acoustic oscillations (BAO) from the Dark Energy Spectroscopic Instrument (DESI) DESI:2024uvr ; DESI:2024lzq , combined with cosmic microwave background (CMB) data Planck:2018vyg and supernova (SN) observations DES:2024tys , it appears DESI:2024mwx that the $\Lambda$CDM model is losing ground, signalling the onset of a dynamical dark energy (DE) era.
@@
The new DESI data DESI:2024mwx have attracted significant interest from the scientific community, prompting the development of various models that align with these observations. These models include quintessence from scalar fields Tada:2024znt ; Berghaus:2024kra ; Shlivko:2024llw ; Bhattacharya:2024hep ; Ramadan:2024kmn , interacting DE Giare:2024smz , dark radiation Allali:2024cji , and DE produced by baryon conversion in cosmologically coupled black holes Croker:2024jfg or in an axion phase transition Muursepp:2024mbb . Modifications of gravity in light of the new DESI data have also been discussed Yang:2024kdo ; Escamilla-Rivera:2024sae . Additionally, numerous scenarios extending beyond the $\Lambda$CDM model have been proposed Yin:2024hba ; Gu:2024jhl ; Wang:2024qan ; Wang:2024hks ; Cortes:2024lgw ; Colgain:2024xqj ; Wang:2024rjd ; Carloni:2024zpl . A thorough analysis of physically motivated DE models (thawing Caldwell:2005tm , emergent Li:2020ybr , mirage class Linder:2007ka ) was presented in Lodha:2024upq . Moreover, the Hubble tension Riess:2016jrr ; Riess:2021jrx in light of the new DESI data has been analyzed in Clifton:2024mdy ; Bousis:2024rnb ; Qu:2024lpx ; Wang:2024dka ; Seto:2024cgo ; Pogosian:2024ykm ; Jia:2024wix ; DES:2024ywx . The validity of distance duality relation using DESI BAO measurements has been investigated in Ref. Favale:2024sdq . A model-independent analysis has been provided by the DESI team Calderon:2024uwn , see also Refs. Luongo:2024fww ; Mukherjee:2024ryz ; Dinda:2024kjf ; DES:2024fdw . The need for evolving DE could also be reached without the new DESI measurements by combining CMB and SN data with older BAO measurements Park:2024jns . Recently, also a mild $2\sigma$ tension was reported between the sum of neutrino masses expected from oscillation experiments and those derived from cosmological data (CMB, SN, DESI BAO) Wang:2024hen .
@@
Our first result is that, when performing global fits to DESI, CMB and SN data, it is important to keep in mind which physical scenario is under study. Using the linear ansatz for $w$, which is motivated by its simplicity but does not represent any physical scenario, gives a misleading impression that the new data prefers phantom DE. The quintessence and step function parametrizations give comparable or better fits to data without any hint of phantom DE. Thus the preference for a $w<-1$ regime is just an artefact of the choice of parametrization.
@@
Modelling dynamical dark energy – We will model DE with a time-dependent equation of state parameter $w(a)$, but with a conserved energy-momentum tensor, that is,
@@
A particularly simple phenomenological parameterization [^1] In light of DESI data DESI:2024mwx , a polynomial parametrization has been used in Yang:2024kdo , while in Escamilla:2024olw , the use of trigonometric functions gives rise to an oscillating equation of state. is given by Chevallier:2000qy ; Linder:2002et
@@
where $w_{0}$ and $w_{a}$ are free parameters that determine the variation of $w$ from $w_{0}$ at the present to $w_{0}+w_{a}$ in the past when $a\to 0$. Due to its simplicity, this parameterization is widely used and was implemented in the recent analysis by DESI collaboration DESI:2024mwx . The dynamics of the DE component is given by
@@
Confronting DESI data – For observational data we use the first year DESI BAO measurements DESI:2024mwx , Planck CMB bounds Planck:2018vyg on angular scale of the sound horizon and supernova luminosity distance measurements from the Dark Energy Survey (DES) Year 5 data release DES:2024tys . DESI provides BAO measurements at seven redshifts: $z=0.30,0.51,0.71,0.93,1.32,1.49,2.33$. For most of the redshifts anisotropic, i.e., along and perpendicular to line-of-sight measurements of BAO, together with their correlation coefficients are provided, while for redshifts $0.30$ and $1.49$ only isotropized BAO measurements are given. Thus, in total $2\times 5+2=12$ BAO measurements, which are assumed to follow Gaussian error distributions, are used in this work. As our analysis does not attempt to achieve absolute calibration of the BAO we work with a combined scaling parameter $r_{d}h$ ($r_{d}$ – sound horizon at the baryon ‘drag’ epoch, $h$ – reduced Hubble parameter). Thus, e.g. the results obtained for the Hubble parameter should be understood as implicitly including an additional factor $r_{d}/r_{d}^{\rm fid}$. We consider only spatially flat models and for our fiducial $\Lambda$CDM model (best-fit Planck 2018 $\Lambda$CDM Planck:2018vyg ) with $\Omega_{m}h^{2}=0.1428$ and $\Omega_{b}h^{2}=0.02233$ sound horizon $r_{d}^{\rm fid}=147.17$ Mpc. Our Markov chain Monte Carlo (MCMC) explorations can be simplified since, e.g. marginalizations – equivalent to maximization under the Gaussianity assumption – over $r_{d}h$, which acts as a common multiplicative scaling parameter, can be done analytically.
@@
Our main results are summarized by Figs. 1-3 and in Table 1. Fig. 1 is devoted to the linear DE model, which has been extensively studied in the literature and as such should serve as a test case for our calculations (corner plots for the quintessence and step function models can be found in the Supplementary Material). The parameter constraint contours ($1\sigma$ and $2\sigma$) are shown for DESI data alone, DESI+CMB, and DESI+CMB+SN data, the last with and without local supernovae, i.e., including or excluding CfA/CSP sample, respectively.
Using the full SN data it is clear that $\Lambda$CDM is disfavored over the linear $w_{0}-w_{a}$-parametrization: as shown in Table 1 the respective $\Delta\chi^{2}$ gets increased by $12.9$. However, if we remove low-$z$ SN, then the analysis does not require evolving DE and thus fully restores $\Lambda$CDM concordance.
@@
As seen from Table 1, the goodness of fit for all three DE models is almost equal, with the step model performing slightly better than the other two.
The $\chi^{2}$ values given in Table 1 might look abnormally low for the number of degrees of freedom we have: $1829$ SNe + $12$ BAO measurements + $1$ CMB constraint -- $4$ model parameters ($\Omega_{m}$, $r_{d}h$, + 2 DE parameters) = 1838. These low values are fully driven by the SN data. As discussed in Ref. DES:2024tys , due to the way contamination of the SN data is treated the effective number of SNe is significantly lower $\sim 1735$, rather than $1829$.
Model $\chi^{2}$ $\Delta\chi^{2}$ $\Omega_{m}$ $w_{0}$ $w_{a}\mid z_{t}\mid K$
Linear $1649.7$ $0.3$ $0.318\pm 0.007$ $-0.735\pm 0.081$ $-1.05\pm 0.42$
$0.318$ $-0.750$ $-0.96$
Step $1649.4$ $0.0$ $0.326\pm 0.009$ $-0.32\pm 0.47$ $0.081\pm 0.048$
@@
Table 1: Goodness of fit, posterior mean and standard deviations for the parametrizations. Best fit values are given in small print.
Fig. 3 shows the evolution of DE density, its equation of state, and the time derivative of the scale factor (the last of which has omitted an implicit factor of $r_{d}/r_{d}^{\rm fid}$). It is evident that the only robust and consistent feature amongst the models is the very recent/local ($z\lesssim 0.1$) sharp drop, ${\cal O}(10\%)$, in the DE density, or equivalently a sharp rise in the DE equation of state parameter. The fast drop in the DE density and corresponding crossing of the phantom barrier at higher redshifts is an artefact of the $w_{0}-w_{a}$-parametrization (quite unavoidable in this setting, since the data demand large negative $w_{a}$), which is not required by the data.
@@
Discussion and summary – The findings of our letter can be summed up by the following three points:
@@
In particular, as illustrated in Fig. 1 (see also Figs. 4 and 5 in the Supplementary Material), the tension between the standard $\Lambda$CDM and the combined data is relieved if the low-$z$ supernova data is omitted.
Contemporary SN samples have become so large that we have entered a regime where the statistical errors can get overwhelmed by systematics. These issues are not easy to handle since SN are complex phenomena whose detailed physics is not well understood. Thus, there are always possibilities for unaccounted-for systematic effects. Also, for nearby SN, the data analysis is more demanding due to the peculiar motions of their host galaxies, which have a significant effect and thus need to be modelled accurately.
@@
The parameter $w_{0}$ remains poorly constrained in the quintessence and the step function parametrizations, with potential values being $w_{0}\gtrsim-0.8$. We find that the preference for a narrow range of $w_{0}$ in the linear parametrization can be considered as another artefact of that ansatz.
@@ Acknowledgements.
Acknowledgments – We would like to thank Ville Vaskonen for the discussions. We acknowledge the use of GetDist software Lewis:2019xzd . This work was supported by the Estonian Research Council grants MOB3JD1202, PSG869, PRG803, PRG1055, PRG1677, PSG761, RVTT3 and RVTT7 and the Center of Excellence program TK202 “Fundamental Universe”.
Supplementary Material
We show here the corresponding corner plots from our likelihood analysis for the quintessence (Fig. 4) and step function (Fig. 5) models.
@@
Figure 5: Parameter constraints ($1\sigma$ and $2\sigma$ contours) for the step function DE model assuming DESI+CMB+SN data with and without local SN, as shown in the legend.

### v2

@@ Abstract
We perform fits to DESI, CMB and supernova data to understand the physical origin of the DESI hint for dynamical dark energy. We find that the parametrization $w=w_{0}+(1-a)w_{a}$ can, misleadingly, hint for a phantom Universe, although quintessence models fit the data well. Model-independently, we find that deviations from $\Lambda$CDM are driven by low-$z$ supernova data and take place only at very low redshifts $z<0.1$. Therefore, either our Universe underwent dramatic changes very recently or we do not fully understand our local Universe in a radius of about $300\,h^{-1}\rm Mpc$.
I Introduction
@@
We have come to a turning point where the standard $\Lambda$CDM model, previously regarded as robust, may have to be reconsidered. Taking into account recent measurements of baryon acoustic oscillations (BAO) from the Dark Energy Spectroscopic Instrument (DESI) DESI:2024uvr ; DESI:2024lzq , combined with cosmic microwave background (CMB) data Planck:2018vyg and supernova (SN) observations DES:2024jxu , it appears DESI:2024mwx that the $\Lambda$CDM model is losing ground, signalling the onset of a dynamical dark energy (DE) era.
@@
The new DESI data DESI:2024mwx have attracted significant interest from the scientific community, prompting the development of various models that align with these observations. These models include quintessence from scalar fields Tada:2024znt ; Berghaus:2024kra ; Shlivko:2024llw ; Bhattacharya:2024hep ; Ramadan:2024kmn , interacting DE Giare:2024smz , dark radiation Allali:2024cji , and DE produced by baryon conversion in cosmologically coupled black holes Croker:2024jfg or in an axion phase transition Muursepp:2024mbb . Modifications of gravity in light of the new DESI data have also been discussed Yang:2024kdo ; Escamilla-Rivera:2024sae . Additionally, numerous scenarios extending beyond the $\Lambda$CDM model have been proposed Yin:2024hba ; Gu:2024jhl ; Wang:2024qan ; Wang:2024hks ; Colgain:2024xqj ; Wang:2024rjd ; Carloni:2024zpl . A thorough analysis of physically motivated DE models (thawing Caldwell:2005tm , emergent Li:2020ybr , mirage class Linder:2007ka ) was presented in Lodha:2024upq . Moreover, the Hubble tension Riess:2016jrr ; Riess:2021jrx in light of the new DESI data has been analyzed in Clifton:2024mdy ; Bousis:2024rnb ; Qu:2024lpx ; Wang:2024dka ; Seto:2024cgo ; Pogosian:2024ykm ; Jia:2024wix ; DES:2024ywx . The validity of distance duality relation using DESI BAO measurements has been investigated in Ref. Favale:2024sdq , separation of the radial and angular BAO measurements with an attempt to isolate the potentially problematic DESI redshift bins was carried out in Ref. Wang:2024pui . A model-independent analysis has been provided by the DESI team DESI:2024aqx , see also Refs. Luongo:2024fww ; Mukherjee:2024ryz ; Dinda:2024kjf ; DES:2024fdw . The need for evolving DE could also be reached without the new DESI measurements by combining CMB and SN data with older BAO measurements Park:2024jns . Similar hints for evolving DE have also been obtained before the DESI BAO measurements Rubin:2023ovl ; DES:2024jxu ; Colgain:2024ksa . Recently, also a mild $2-3\sigma$ tension was reported between the sum of neutrino masses expected from oscillation experiments and those derived from cosmological data (CMB, SN, DESI BAO) Craig:2024tky ; Wang:2024hen .
@@
Our first result is that, when performing global fits to DESI, CMB and SN data, it is important to keep in mind which physical scenario is under study. Using the linear ansatz for $w$, which is motivated by its simplicity but does not represent any physical scenario, gives a misleading impression that the new data prefers phantom DE. Phantom models can be problematic because they violate fundamental principles, such as the null energy condition, and lead to unstable vacuum states due to unbounded Hamiltonians, rendering them incompatible with well-defined particle physics theories. The quintessence and step function parametrizations give comparable or better fits to data without any hint of phantom DE. Thus the preference for a $w<-1$ regime is just an artefact of the choice of parametrization.
@@
II Modelling dynamical dark energy
We will model DE with a time-dependent equation of state parameter $w(a)$, but with a conserved energy-momentum tensor, that is,
@@ II Modelling dynamical dark energy
A particularly simple phenomenological parametrization [^1] In light of DESI data DESI:2024mwx , a polynomial parametrization has been used in Yang:2024kdo , while in Escamilla:2024fzq , the use of trigonometric functions gives rise to an oscillating equation of state. is given by Chevallier:2000qy ; Linder:2002et
@@
where $w_{0}$ and $w_{a}$ are free parameters that determine the variation of $w$ from $w_{0}$ at the present to $w_{0}+w_{a}$ in the past when $a\to 0$. Due to its simplicity, this parametrization is widely used and was implemented in the recent analysis by DESI collaboration DESI:2024mwx (for an extended discussion of this parametrization and its shortcomings see e.g. Wolf:2023uno ; Cortes:2024lgw ). The dynamics of the DE component is given by
@@
III Confronting DESI data
For observational data we use the first year DESI BAO measurements DESI:2024mwx , Planck CMB bounds Planck:2018vyg on angular scale of the sound horizon and supernova luminosity distance measurements from the Dark Energy Survey (DES) Year 5 data release DES:2024jxu . DESI provides BAO measurements at seven redshifts: $z=0.30,0.51,0.71,0.93,1.32,1.49,2.33$. For most of the redshifts anisotropic, i.e., along and perpendicular to line-of-sight measurements of BAO, together with their correlation coefficients are provided, while for redshifts $0.30$ and $1.49$ only isotropized BAO measurements are given. Thus, in total $2\times 5+2=12$ BAO measurements, which are assumed to follow Gaussian error distributions, are used in this work.  [^2] In our analysis we also implicitly assume that the standard BAO analysis as carried out by the DESI collaboration (which as an input needs assumptions about the fiducial cosmology) is sufficient for our purposes. In case real cosmology deviates significantly from the assumed fiducial model one should rather rely on ‘purely-geometric BAO’ methods as suggested in ODwyer:2019rvi ; Anselmi:2022exn . Additionally, it is important to note that this approximate analysis tends to underestimate the statistical uncertainties. As our analysis does not attempt to achieve absolute calibration of the BAO we work with a combined scaling parameter $r_{d}h$ ($r_{d}$ – sound horizon at the baryon ‘drag’ epoch, $h$ – reduced Hubble parameter). Thus, e.g. the results obtained for the Hubble parameter should be understood as implicitly including an additional factor $r_{d}/r_{d}^{\rm fid}$. We consider only spatially flat models [^3] Including spatial curvature is beyond the scope of this study. However, a similar analysis is possible for non-flat geometries ($\Omega_{K}\neq 0$Ichikawa:2005nb ; Clarkson:2007bc ; Anselmi:2022uvj . and for our fiducial $\Lambda$CDM model (best-fit Planck 2018 $\Lambda$CDM Planck:2018vyg ) with $\Omega_{m}h^{2}=0.1428$ and $\Omega_{b}h^{2}=0.02233$ sound horizon $r_{d}^{\rm fid}=147.17$ Mpc. Our Markov chain Monte Carlo (MCMC) explorations can be simplified since, e.g. marginalizations – equivalent to maximization under the Gaussianity assumption – over $r_{d}h$, which acts as a common multiplicative scaling parameter, can be done analytically.
@@ III Confronting DESI data
Our main results are summarized by Figs. 1-3 and in Table 1. Fig. 1 is devoted to the linear DE model, which has been extensively studied in the literature and as such should serve as a test case for our calculations (corner plots for the quintessence and step function models can be found in the Appendix). The parameter constraint contours ($1\sigma$ and $2\sigma$) are shown for DESI data alone, DESI+CMB, and DESI+CMB+SN data, the last with and without local supernovae, i.e., including or excluding CfA/CSP sample, respectively.
From Fig. 1 it is evident that to achieve a reasonable measurement for the DE parameters one needs the input from all the datasets, e.g. CMB+DESI data without SN provides only a bound on the linear combination of $w_{0}$ and $w_{a}$ (with a best fit, even though this is not particularly meaningful, away from the $\Lambda$CDM point). Similar strongly degenerate measurements, albeit already showing some hints for mild deviations from $\Lambda$CDM, were obtained earlier using SN+CMB data, see e.g. DES:2024jxu .
Using the full SN data it is clear that $\Lambda$CDM is disfavored over the linear $w_{0}-w_{a}$-parametrization: as shown in Table 1 the respective $\Delta\chi^{2}$ gets increased by $12.9$, or $\Delta$AIC 1974ITAC…19..716A by 8.9. However, if we remove low-$z$ SN, then the analysis does not require evolving DE and thus fully restores $\Lambda$CDM concordance, the fact most clearly demonstrated by the $\Delta\chi^{2}$ and $\Delta$AIC values shown in Table 2. Along similar lines, without the low-$z$ SN, we find that $z_{t}$ (and $K$) maintain an upper (lower) bound, but the lower (upper) bound, that was present with the low-$z$ SN, vanishes and $\Lambda$CDM (that is, $z_{t}=0$ and $K\to\infty$) is now accommodated by the data. Thus it follows that the preference for DE over $\Lambda$CDM is driven by the low-$z$ SN data.
@@
As seen from Table 1, the goodness of fit for all three DE models is almost equal, with the step model performing slightly better than the other two. The preference of evolving DE models over $\Lambda$CDM gets somewhat reduced once penalizing their extra degree of complexity, which according to the Akaike information criterion 1974ITAC…19..716A leads to the $\Delta$AIC values as shown in the table.
Model $\chi^{2}$ $\Delta\chi^{2}$ $\Delta$AIC $\Omega_{m}$ $w_{0}$ $w_{a}\mid z_{t}\mid K$
Linear $1649.7$ $0.3$ $0.3$ [0.31, 0.33] [-0.89, -0.57] [-1.92,-0.26]
$0.318$ $-0.75$ $-0.96$
Step $1649.4$ $0.0$ $0.0$ [0.31, 0.34] [-0.93, 0.72] [0.02,0.19]
@@
Table 1: Goodness of fit and the 95.45% confidence intervals for the different parametrizations. Best fit values are given in small print below.
Model $\chi^{2}$ $\Delta\chi^{2}$ $\Delta$AIC $\Omega_{m}$ $w_{0}$ $w_{a}\mid z_{t}\mid K$
Linear $1465.5$ $0.0$ $3.1$ [0.29, 0.33] [-1.14, -0.57] [-1.64, 0.54]
$0.306$ $-0.90$ $-0.38$
Step $1466.3$ $0.8$ $3.9$ [0.30, 0.36] $z_{t}<0.18$
$0.315$ $-0.66$ $0.060$
Quintessence $1466.3$ $0.8$ $3.9$ [0.30, 0.34] $K>7.7$
$0.320$ $0.56$ $24.5$
$\Lambda$CDM $1466.4$ $0.9$ $0.0$ [0.29, 0.32]
$0.304$
Table 2: Same as Table 1 but for the analysis without the local SN data. The $w_{0}$ confidence intervals are not shown when the posterior is essentially flat.
The $\chi^{2}$ values given in Table 1 might look abnormally low for the number of degrees of freedom we have: $1829$ SNe + $12$ BAO measurements + $1$ CMB constraint -- $4$ model parameters ($\Omega_{m}$, $r_{d}h$, + 2 DE parameters) = 1838. These low values are fully driven by the SN data. As discussed in Ref. DES:2024jxu , due to the way contamination of the SN data is treated the effective number of SNe is significantly lower $\sim 1735$, rather than $1829$.
Fig. 3 shows the evolution of DE density, its equation of state, and the time derivative of the scale factor (the last of which has omitted an implicit factor of $r_{d}/r_{d}^{\rm fid}$). It is evident that the only robust and consistent feature amongst the models is the very recent/local ($z\lesssim 0.1$) sharp drop, ${\cal O}(10\%)$, in the DE density (for similar findings on this topic, see Shlivko:2024llw ), or equivalently a sharp rise in the DE equation of state parameter. The fast drop in the DE density and corresponding crossing of the phantom barrier at higher redshifts is an artefact of the $w_{0}-w_{a}$-parametrization (quite unavoidable in this setting, since the data demand large negative $w_{a}$), which is not required by the data.
@@
It is instructive to compare our results with a “model agnostic” study DESI:2024aqx , which uses a relatively low order Chebyshev polynomial basis for $w(a)$. Together with a sharp rise at low $z$, they have a mild preference for phantom behaviour at higher redshifts. Polynomials, however, do not naturally permit rapid changes and thus the high-$z$ behaviour can be dictated by the low-$z$ one, if the latter contains sudden changes in $w$.
IV Discussion and conclusions
The findings of our paper can be summed up by the following three points:
@@ IV Discussion and conclusions
In particular, as illustrated in Fig. 1 (see also Figs. 4 and 5 in the Appendix), the tension between the standard $\Lambda$CDM and the combined data is relieved if the low-$z$ supernova data is omitted.
Contemporary SN samples have become so large that we have entered a regime where the statistical errors can get overwhelmed by systematics. These issues are not easy to handle since SN are complex phenomena whose detailed physics is not well understood. Thus, there are always possibilities for unaccounted-for systematic effects. Also, for nearby SN, the data analysis is more demanding due to the peculiar motions of their host galaxies, which have a significant effect and thus need to be modeled accurately. [^6] Although our analysis including low-$z$ SN data suggests that deviation from standard $\Lambda$CDM takes place only at low $z$, we cannot rule out that the tension with $\Lambda$CDM might be relieved by considering other data splits that involve excluding data from higher redshifts because we have not considered such possibilities here.
@@
The parameter $w_{0}$ remains poorly constrained in the quintessence and the step function parametrizations, with potential values being $w_{0}\gtrsim-0.6$ and $w_{0}\gtrsim-0.9$ respectively. The preference for a narrow range of $w_{0}$ in the linear parametrization can be considered another artefact of that ansatz.
@@ Acknowledgements.
Acknowledgments – We would like to thank Ville Vaskonen for the discussions. We acknowledge the use of GetDist software Lewis:2019xzd . This work was supported by the Estonian Research Council grants MOB3JD1202, PSG869, PRG803, PRG1055, PRG1677, RVTT3 and RVTT7 and the Center of Excellence program TK202 “Fundamental Universe”.
Note added – During the refereeing process of this paper, the work Efstathiou:2024xcq appeared whose findings regarding the impact of local SN are in line with our results.
Appendix A Appendix: Corner plots for DE fits
We show here the corresponding corner plots from our likelihood analysis for the quintessence (Fig. 4) and step (Fig. 5) models.
@@ Appendix A Appendix: Corner plots for DE fits
Figure 5: Parameter constraints ($1\sigma$ and $2\sigma$ contours) for the step DE model assuming DESI+CMB+SN data with and without local SN, as shown in the legend.